OpenStudy (anonymous):

3/8 divided by -2 2/3 = ???

OpenStudy (skullpatrol):

$$\huge \frac{3}{8}\div (-2\frac{2}{3})$$

OpenStudy (skullpatrol):

The first thing you need to do is convert the mixed number into a fraction, can you do that?

OpenStudy (skullpatrol):

$$\huge-2\frac{2}{3} =?$$

OpenStudy (anonymous):

Uhhh remind me how to do that again...

OpenStudy (skullpatrol):

$$ 2\frac{2}{3}$$$$=2+\frac{2}{3}$$$$=\frac{6}{3}+\frac{2}{3}$$$$=\frac{6+2}{3}$$$$=\frac{8}{3}$$

OpenStudy (anonymous):

oh thanks!

OpenStudy (skullpatrol):

Now you can use the definition of division: $$\huge \frac{a}{b}\div \frac{c}{d} =\frac{a}{b}*\frac{d}{c}$$

OpenStudy (skullpatrol):

$$\huge \frac{3}{8}\div (-\frac{8}{3})=\huge \frac{3}{8}* (-\frac{3}{8})$$

OpenStudy (skullpatrol):

To multiply two fractions, you multiply their numerators and multiply their denominators.